# Communication-Efficient Wireless Federated Fine-Tuning for Large-Scale AI Models

### v1

@@ III-A Loss Function Design
A fundamental component of our sparsification method is the design of a loss function that enforces orthogonal properties on the matrices $\theta_{B,k}$ and $\theta_{A,k}$. This design is inspired by the orthogonality inherent in the singular matrices of SVD. Specifically, we aim for $\theta_{B,k}^{\textsf{T}}\theta_{B,k}$ and $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ to approximate diagonal matrices, analogous to the roles of $U$ and $V$ in SVD. To accomplish this, each client $k$ constructs the following loss function:
@@
where $L_{task,k}$ denotes the task-specific loss for client $k$, $\zeta$ is a scaling parameter, and $\text{diag}(C)$ denotes a diagonal matrix with the diagonal elements of $C$.
Fig. 2 illustrates the impact of our proposed loss function on promoting orthogonality within $\theta_{B,k}$ and $\theta_{A,k}$. After training with this loss function, the product $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ becomes approximately diagonal, demonstrating the successful enforcement of orthogonal properties. This property allows us to compute the Frobenius norm by aggregating the $\ell_{2}$-norms of the individual vectors across the $r$ components, as expressed mathematically $\|\theta_{B,k}\theta_{A,k}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2},$ where $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$ are the $i$-th column and row vectors, respectively, whose orthogonality ensures that cross terms vanish, simplifying the calculation to sum of vector norms.
Figure 2: Visualization of $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ of Vit-Base blocks.6.mlp.fc1 layer: without (a) and with (b) the proposed loss function.
@@ III-B Sparsification Method
Since $\theta_{B,k}$ and $\theta_{A,k}$ exhibit orthogonal properties, we can compute singular values by calculating the norms of the matrices $\theta_{B,k}$ and $\theta_{A,k}$. For instance, the $i$-th singular value, where $i=\{1,...,r\}$, can be obatained as $\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}$. Because the magnitude of singular values represents the importance of their corresponding singular vectors, we implement sparsification based on the magnitude of these singular values. For each singular vector, we define $o_{k,i}$ as the number of non-zero elements retained in both $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. The value of $o_{k,i}$ is determined by distributing the total number of non-sparse elements proportionally to the magnitude of the singular values. Specifically, $o_{k,i}$ is calculated as follows:
@@
Finally, based on (3), client $k$ selects top-$o_{k,i}$ elements in $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. Although the proposed sparsification method also requires calculating the $l$-2 norm $r$ times in each iteration, this computation has a complexity of $\mathcal{O}(r(d+\ell))$, negligible compared to the original complexity caused by full matrix multiplication and SVD operation, i.e. $\mathcal{O}(dr\ell+\min(d,\ell)\cdot d\ell)$.
Algorithm 1 Sparsified Orthogonal Fine-Tuning
1:Input: The global model $\theta_{B}$ and $\theta_{A}$, sparsification ratio $O_{k}$.
2:for $k\in\mathcal{K}^{t}$ $\triangleright$ Local update in parallel
3:    Download $\theta_{B_{k}}^{t,0}=\theta_{B}$ and $\theta_{A_{k}}^{t,0}=\theta_{A}$ from the PS
4:    for $e=0,\dots,E-1$ do
5:       Locally iterate using (III-A)
6:    end for
7:    Incorporate error feedback:
8:    $m_{B,k}^{t}+\Delta\theta_{B,k}^{t},m_{A,k}^{t}+\Delta\theta_{A,k}^{t}$
9:    Compute sparified updates $\theta_{B,k}$ and $\theta_{A,k}$ with rate $O_{k}$
10:    Transmit $\theta_{B,k}^{t}$ and $\theta_{A,k}^{t}$ to the server
11:end for
@@ III-C Error Feedback with SOFT
While the proposed sparsification method effectively reduces communication costs by transmitting only the most significant elements of $\theta_{B,k}$ and $\theta_{A,k}$, it inevitably introduces sparsification errors. To mitigate the impact of these errors on model convergence and performance, we incorporate an error feedback mechanism into SOFT. The error feedback mechanism works by accumulating the sparsification errors at each client and incorporating them into the updates [9]. Specifically, each client maintains local error memory matrices, denoted as $m_{B,k}$ and $m_{A,k}$, which store the cumulative sparsification errors for $\theta_{B,k}$ and $\theta_{A,k}$, respectively. The sparified update can be expressed as $\theta_{B,k}^{t}=\mathcal{S}(m_{B,k}^{t}+\Delta\theta_{B,k}^{t}),\theta_{A,k}^{t}=\mathcal{S}(m_{A,k}^{t}+\Delta\theta_{A,k}^{t}),$ where $\mathcal{S}$ operates the sparsification algorithm, $\Delta\theta_{B,k}^{t}$ and $\Delta\theta_{A,k}^{t}$ represent obtained model of client $k$ at time $t$ using the loss function (III-A). Subsequently, the error memory is updated to accumulate the residuals resulting from sparsification:
@@
$\Delta\tilde{\theta}_{k}^{t}\!=\!\begin{bmatrix}{\Delta\theta_{B,k}^{t\textsf{T}}}\Delta\theta_{A,k}^{t}\end{bmatrix},\tilde{\theta}_{k}^{t}\!=\!\begin{bmatrix}{\theta_{B,k}^{t\textsf{T}}}\theta_{A,k}^{t}\end{bmatrix},$ (6)
where []delimited-[][$\cdot$][ ⋅ ] is concatenation operation. With these definitions, the overall error memory update can be succinctly written as
@@
This process effectively corrects the errors introduced by sparsification, ensuring that the model updates are more accurate over time. The error feedback mechanism enhances the convergence of the FL process despite the sparsification. The advantage of our proposed sparsification method will be evaluated in Subsection V-A and summarized in Algorithm 1.
We state a lemma that shows the errors maintained in Algorithm 1 using the following assumption[9].
@@
Under the orthogonal properties, sparsification rate $O_{k}$ and Assumption 1, the expected squared norm of the sparsification error $\tilde{m}_{k}^{t+1}$ at iteration $t$ is upper-bounded as
@@
Invoking Lemma 1, we complete the proof. ∎
@@ IV-A Convergence Analysis and Problem Formulation
In LoRA-based FL, unlike conventional FL approaches, the use of LoRA modules introduces two additional factors that must be considered in the learning process. First, as discussed in Subsection II-A, the performance of LoRA is predominantly influenced by the rank $r$. Therefore, an analysis of the impact of the selected rank for fine-tuning should be included. To quantify the original fully-finetuning model and LoRA model with rank $r$, we introduce the following assumption including an inequality.
@@
Let $\theta_{o}$ denote the LoRA model that achieves the same performance as the fully fine-tuned model, and let $\theta_{r}$ be the LoRA model of rank $r$ trained using the loss function (III-A). Denote $\sigma(\theta_{r})$ the largest singular value of $\theta_{r}$, i.e. $\sigma(\theta_{r})=\max_{i}\|\theta_{B}[:,i]\|_{2}\|\theta_{A}[i,:]\|_{2}$, for $i=1,\dots,r$. Then, there exists a constant $H>0$ such that the LoRA error satisfies:
@@
Second, as mentioned in Subsection II-B, clients transmit the matrices $\theta_{B}$ and $\theta_{A}$ instead of the full matrix product $\theta_{B}\theta_{A}$. While this approach reduces computational overhead, it introduces an unexpected additional term when aggregating the updates at the server. The additional term introduced in the aggregation at the sever is as follows.
@@
$=\mathbb{E}_{k}[(\theta_{B,k}-\theta_{B})](\theta_{A,k}-\theta_{A})]=\textit{Cov}(\theta_{B},\theta_{A}).$ (12)
@@
To formally account for this behavior, we introduce the following assumption:
@@
Let $\theta_{r}$ denote the LoRA model with rank $r$. The expected Frobenius norm of the covariance of $\theta_{B}$ and $\theta_{A}$ satisfies:
$\mathbb{E}[\|\textit{Cov}(\theta_{B},\theta_{A})\|_{F}^{2}]\leq\phi\|\theta_{r}\|_{F}^{2},$ (13)
where $\phi$ is a constant that quantifies the degree of data heterogeneity.
@@
Assumption 3 indicates that the covariance between $\theta_{B}$ and $\theta_{A}$ is significantly affected by both the diversity of client data and the size of the model. Empirical evidence supporting this relationship is provided in Subsection V-C.
@@
Building upon the convergence analysis , our goal is to jointly optimize the key parameters to minimize the convergence upper bound. Specifically, we aim to determine the optimal LoRA rank $r$, sparsification ratios $O^{t}$, and bandwidth allocations $b_{k}$ to enhance both communication efficiency and learning performance under the constraints of limited bandwidth and training time. Our optimization problem is formulated as follows.
@@
where $\gamma^{T}\!=\!\frac{1}{T}\!\sum_{t\!=\!0}^{T\!-\!1}2S^{2}H(r\!_{\max}\!-\!r)\!W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}+\frac{8(N\!-\!K^{t})}{K^{t}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1\!-\!O^{t})^{2}}{K^{t}{O^{t}}^{4}}rS^{2}W^{4}+\frac{E(E\!-\!1)(2E\!-\!1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the optimality gap in Theorem 1; (15b) is the rank feasibility constraint; (15c) is the average transmission delay constraint; (15d) is the sparsification ratio constraint with minimum sparsification ratio $O_{min}$; (15e) is the constraints for the bandwidth allocation strategy.
Note that Problem $\mathcal{P}_{1}$ presents two key challenges. First, it must be addressed prior to training, as the model structure, i.e. LoRA rank $r$, needs to be determined in advance. However, this approach would require multiple training iterations and full knowledge of the channel state information (CSI) for all clients at each iteration, which is impractical in the real scenario. Second, the presence of long-term constraints, highly coupled variables, and mixed-integer non-linear programming (MINLP) problem further makes the problem NP-hard. To address these challenges, we divide the problem into two stages. In the offline stage, we optimize the LoRA rank $r$ before training begins by analyzing the trade-off between model capacity and communication overhead. In the online stage, we adaptively manage the exact parameters except $r$ during training, based on real-time network conditions and resource availability.
@@ IV-B Offline Stage
As discussed in the previous subsection, the offline stage is conducted prior to training with the specific goal of optimizing the appropriate LoRA rank $r$. Since the actual CSI cannot be obtained during the offline stage, the channel is approximated by assuming the average channel as the actual channel. The initial sparsification ratio are set to $O^{0}$. The offline stage thus considers the following approximated problem:
@@
where $\gamma^{0}=2S^{2}H(r_{\max}\!-\!r)W^{2}\!+\!4S^{2}\phi rW^{2}\!+\!\eta SE^{2}G^{2}\!+\!\frac{8(N\!-\!K^{0})}{K^{0}(N\!-\!1)}\eta^{2}S^{2}E^{2}G^{2}\!+\!\frac{4N(1-O^{0})^{2}}{K^{0}{O^{0}}^{4}}rS^{2}W^{4}\!+\!\frac{E(E-1)(2E-1)S^{2}\eta^{2}}{6}G^{2}$, which is the upper-bound on the initial optimality gap in Theorem 1 when $T=0$.
Note that the Problem $\mathcal{P}_{2}$ is still an NP-hard MINLP problem due to the discrete rank value and non-convexity caused by the coupled variables. To address these challenges, we first fix the integer variable $r$ to a specific value and solve the resulting continuous optimization problem $O^{0}$. By iterating over all feasible values of $r$ within the constraint, we can identify the optimal integer $r$. The overall process is in the Algorithm 2. To efficiently solve the continuous optimization problem for a fixed $r$, we present the following lemma, which provides the optimal values of $O^{0}$ given an integer $r$.
Algorithm 2 Offline Stage Optimization Algorithm
1:Initialize $\gamma^{0}_{\text{opt}}\leftarrow\infty$
2:for $r=1$ to $r_{\text{max}}$ do
3:    Fix the LoRA rank $r$
4:    Solve for $O^{0}$ using Lemma 2
5:    Compute the objective function $\gamma^{0}$ with $r$ and $O^{0}$
6:    if $\gamma^{0}<\gamma^{0}_{\text{opt}}$ then
7:       Update $\bar{r}\leftarrow r$
8:    end if
9:end for
10:return $\bar{r}$
@@
$O^{0}=\max\left(O_{min},\min\left(1,\frac{N}{AK^{0}r}\right)\right),$ (17)
@@
Consequently, using the property $O_{min}<O^{0}\leq 1$, we can obtain
$O^{0}=\max\left(O_{min},\min\left(1,\frac{N}{AK^{0}r}\right)\right),$ (20)
@@
Based on Lemma 2 and Algorithm 2, we can obtain the optimal LoRA rank to be used in training. Subsequently, clients perform fine-tuning based on the specified LoRA rank $r$. At each iteration, an online-stage optimization is conducted to determine the optimal values of $O^{t}$ and $b_{k}^{t}$.

### v2

@@ III-A Loss Function Design
To reduce communication overhead in FL, model updates are commonly sparsified before transmission. An effective sparsification rule should therefore preserve the most informative components of the update. Among rank-$r$ representations of a matrix, truncated SVD provides a desirable approximation in the Frobenius norm [21], suggesting that dominant singular directions offer a principled criterion for importance. However, computing SVD at every communication round is prohibitively expensive for resource-constrained clients in LoRA-based FL. We therefore seek a lightweight surrogate that captures this structural property without explicit decomposition.
A fundamental component of our sparsification method is the design of a loss function that enforces nearly orthogonal properties on the matrices $\theta_{B,k}$ and $\theta_{A,k}$. Specifically, we aim for $\theta_{B,k}^{\textsf{T}}\theta_{B,k}$ and $\theta_{A,k}\theta_{A,k}^{\textsf{T}}$ to approximate diagonal matrices, analogous to the roles of $U$ and $V$ in SVD. Achieving this ensures that the column vectors of $\theta_{B,k}$ and the row vectors of $\theta_{A,k}$ are nearly orthogonal, thereby simplifying subsequent computations. To accomplish this, each client $k$ constructs the following loss function:
@@
where $L_{task,k}$ denotes the task-specific loss for client $k$, $\zeta$ is a scaling parameter, and $\text{diag}(C)$ denotes a diagonal matrix with the diagonal elements of $C$. This loss function ensures that $\theta_{B,k}$ and $\theta_{A,k}$ approximate the orthogonal properties of the $U$ and $V$ matrices in SVD, respectively. The effect of the proposed loss function and the resulting performance are examined in detail in Subsection V-D. Therefore, this property allows us to compute the Frobenius norm by aggregating the $\ell_{2}$-norms of the individual vectors across the $r$ components, as expressed mathematically
$\|\theta_{B,k}\theta_{A,k}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2},$ (8)
where $\theta_{B,k}[:,i]\!$ and $\theta_{A,k}[i,:]\!$ are the $i$-th column and row vectors, respectively, whose near orthogonality ensures that cross terms vanish, simplifying the calculation to sum of vector norms.
@@ III-B Sparsification Method
Since $\theta_{B,k}$ and $\theta_{A,k}$ exhibit nearly orthogonal properties, we can compute singular values by calculating the norms of the matrices $\theta_{B,k}$ and $\theta_{A,k}$. For instance, the $i$-th singular value, where $i=\{1,...,r\}$, can be obtained as $\|\theta_{B,k}[:,i]\|_{2}^{2}\|\theta_{A,k}[i,:]\|_{2}^{2}$. Because the magnitude of singular values represents the importance of their corresponding singular vectors, we implement sparsification based on the magnitude of these singular values. For each singular vector, we define $o_{k,i}$ as the number of non-zero elements retained in both $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. The value of $o_{k,i}$ is determined by distributing the total number of non-sparse elements proportionally to the magnitude of the singular values. Specifically, $o_{k,i}$ is calculated as follows:
@@
Figure 2: Computational complexity comparison between SOFT and the conventional full-multiplication and SVD-based approach. (a) Complexity versus the LoRA rank $r$. (b) Complexity versus matrix dimension.
Finally, based on (9), client $k$ selects top-$o_{k,i}$ elements in $\theta_{B,k}[:,i]$ and $\theta_{A,k}[i,:]$. Although the proposed sparsification method also requires calculating the ll-2 norm $r$ times in each iteration, this computation has a complexity of $\mathcal{O}(r(d+\ell))$, negligible compared to the original complexity caused by full matrix multiplication and SVD operation, i.e. $\mathcal{O}(dr\ell+\min(d,\ell)\cdot d\ell)$. To further illustrate the computational efficiency of SOFT, Fig. 2 compares its computational complexity with that of the conventional full-multiplication and SVD-based approach under varying LoRA ranks and matrix dimensions. In Fig. 2, we fix the matrix dimensions as $d=\ell=4096$ and vary the LoRA rank $r$, whereas in Fig. 2, we fix the LoRA rank as $r=8$ and vary the matrix dimension with $d=\ell$. The $y$-axis is plotted on a logarithmic scale due to the large gap in complexity between the two approaches. This result confirms that the additional overhead of the proposed sparsification method is negligible compared to the original computation.
Remark 1.
Since the loss function (III-A) enforces $\theta_{B}^{\textsf{T}}\theta_{B}$ and $\theta_{A}\theta_{A}^{\textsf{T}}$ to be approximately diagonal, the cross terms in $\|\theta_{B}\theta_{A}\|_{F}^{2}$ vanish and it holds $\|\theta_{B}\theta_{A}\|_{F}^{2}=\sum_{i=1}^{r}\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2},$ so the per-rank score $\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2}$ acts as a surrogate singular value. Performing the sparsification with the per-rank non-zeros $o_{k,i}\propto\|\theta_{B}[:,i]\|_{2}^{2}\|\theta_{A}[i,:]\|_{2}^{2}$ is therefore a consistent SVD-proxy selection rule. Although the loss function (III-A) can slightly reduce accuracy when used alone by encouraging near orthogonality, incorporating the subsequent sparsification step improves overall performance and thus produces favorable accuracy–efficiency trade-offs.
@@ III-C Error Feedback with SOFT
While the proposed sparsification method effectively reduces communication costs by transmitting only the most significant elements of $\theta_{B,k}$ and $\theta_{A,k}$, it inevitably introduces sparsification errors. To mitigate the impact of these errors on model convergence and performance, we incorporate an error feedback mechanism into SOFT. The error feedback mechanism works by accumulating the sparsification errors at each client and incorporating them into the updates [8]. Specifically, each client maintains local error memory matrices, denoted as $m_{B,k}$ and $m_{A,k}$, which store the cumulative sparsification errors for $\theta_{B,k}$ and $\theta_{A,k}$, respectively. The sparsified update can be expressed as
$\theta_{B,k}^{t}=\mathcal{S}(m_{B,k}^{t}+\Delta\theta_{B,k}^{t}),\theta_{A,k}^{t}=\mathcal{S}(m_{A,k}^{t}+\Delta\theta_{A,k}^{t}),$ (10)
where $\mathcal{S}$ operates the sparsification algorithm, $\Delta\theta_{B,k}^{t}$ and $\Delta\theta_{A,k}^{t}$ represent obtained model of client $k$ at time $t$ using the loss function (III-A). Subsequently, the error memory is updated to accumulate the residuals resulting from sparsification:
@@
$\Delta\tilde{\theta}_{k}^{t}=\begin{bmatrix}{\Delta\theta_{B,k}^{t\textsf{T}}}\Delta\theta_{A,k}^{t}\end{bmatrix},\quad\tilde{\theta}_{k}^{t}=\begin{bmatrix}{\theta_{B,k}^{t\textsf{T}}}\theta_{A,k}^{t}\end{bmatrix}$, (13)
where $\cdot$] is concatenation operation. With these definitions, the overall error memory update can be succinctly written as
@@
This process effectively corrects the errors introduced by sparsification, ensuring that the model updates are more accurate over time. The error feedback mechanism enhances the convergence of the FL process despite the sparsification. The advantage of SOFT will be evaluated in Subsection V-A.
We now state a lemma that shows the errors caused by the SOFT using the following assumption[8].
@@
Under nearly orthogonal properties, sparsification rate $O_{k}$ and Assumption 1, the expected squared norm of the sparsification error $\tilde{m}_{k}^{t+1}$ at iteration $t$ is upper-bounded as
@@
Invoking Lemma 1, we have
@@ IV-A Convergence Analysis and Problem Formulation
In LoRA-based FL, unlike conventional FL approaches, the use of LoRA modules introduces two additional factors that must be considered in the learning process. First, as discussed in Subsection II-A, the performance of LoRA is predominantly influenced by the rank $r$. Therefore, an analysis of the impact of the selected rank for fine-tuning should be included. To relate the fully fine-tuned update to the rank $r$ LoRA update, following [22], we introduce the following assumption in the form of an inequality.
@@
Let $\theta_{o}$ denote the LoRA model that achieves the same performance as the fully fine-tuned model, and let $\theta_{r}$ be the LoRA model of rank $r$ trained using the loss function (III-A). Denote $\sigma(\theta_{r})$ as a proxy for the dominant scale of the rank-$r$ LoRA update, i.e. $\sigma(\theta_{r})=\max_{i}\mathbb{E}\|\theta_{B}[:,i]\|_{2}\|\theta_{A}[i,:]\|_{2}$, for $i=1,\dots,r$. Then, there exists a constant $H>0$ such that the LoRA error satisfies:
@@
Second, as mentioned in Subsection II-B, clients transmit the matrices $\theta_{B}$ and $\theta_{A}$ instead of the full matrix product $\theta_{B}\theta_{A}$. While this approach reduces computational overhead, it introduces an unexpected additional term when aggregating the updates at the server [17, 23]. Specifically, since the server averages $\theta_{A,k}$ and $\theta_{B,k}$ independently, it introduce mathematical errors to the global model, i.e., $\mathbb{E}_{k}[\theta_{B,k}\theta_{A,k}]\neq\mathbb{E}_{k}[\theta_{B,k}]\mathbb{E}_{k}[\theta_{A,k}]$. The additional noise term introduced in the aggregation at the server can be expressed by the covariance term $Cov(\cdot)$ as follows.
@@
$=\mathbb{E}_{k}[(\theta_{B,k}-\theta_{B})(\theta_{A,k}-\theta_{A})]$ (20)
$=\sum_{i=1}^{r}\mathbb{E}\left[\left(\theta_{B,k}[:,i]-\theta_{B}[:,i]\right)\left(\theta_{A,k}[i,:]-\theta_{A}[i,:]\right)\right]$ (21)
$=\textit{Cov}(\theta_{B},\theta_{A}).$ (22)
@@
To formally account for this behavior, we introduce the following proposition:
@@
Let $\theta_{r}$ denote the LoRA model with rank $r$. The expected Frobenius norm of the covariance of $\theta_{B}$ and $\theta_{A}$ satisfies $\mathbb{E}[\|\textit{Cov}(\theta_{B},\theta_{A})\|_{F}^{2}]\leq\phi\|\theta_{r}\|_{F}^{2},$ where $\phi$ is a constant that quantifies the degree of data heterogeneity across the clients.
@@
Assumption 3 indicates that, in the LoRA-based FL framework, the covariance between $\theta_{B}$ and $\theta_{A}$ is significantly affected by both the diversity of client data and the size of the model. Empirical evidence supporting this relationship is provided in Subsection V-D.
@@
Building upon the convergence analysis presented in Theorem 1, our goal is to jointly optimize the key parameters to minimize the convergence upper bound. Specifically, we aim to determine the optimal LoRA rank $r$, sparsification ratios $O^{t}$, and bandwidth allocations $b_{k}$ to enhance both communication efficiency and learning performance under the constraints of limited bandwidth and training time. Our optimization problem is formulated as follows.
@@
where (24b) is the rank feasibility constraint; (24c) is the average transmission delay constraint  [^1] In this paper, the delay constraint denotes the latency to upload a client’s model to the server in each iteration.; (24d) is the sparsification ratio constraint with minimum sparsification ratio $O_{min}$; (24e) is the constraints for the bandwidth allocation strategy.
Problem $\mathcal{P}_{1}$ poses two main challenges. First, the structural LoRA rank $r$ must be determined before training because it fixes the model structure and directly affects the model performance as well as the communication payload. Therefore, unlike $O_{k}^{t}$ and $b_{k}^{t}$, $r$ is not an online control variable. Although some prior works [16, 17, 18] consider adaptive rank strategies, they still operate within a pre-defined structural rank budget and do not resolve the structural rank selection problem considered here. Moreover, selecting $r$ online in wireless FL would require repeated evaluation of candidate ranks under time-varying CSI, leading to prohibitive training and signaling overhead. Second, $\mathcal{P}_{1}$ is a MINLP with long-term latency constraints and strongly coupled variables, which makes direct online optimization intractable in practice. We therefore adopt a two-stage design: $r$ is selected offline before training, while $O_{k}^{t}$ and $b_{k}^{t}$ are optimized online during training.
@@ IV-B Offline Stage
As discussed in the previous subsection, the offline stage is conducted prior to training with the specific goal of optimizing the appropriate LoRA rank $r$. Since the actual CSI cannot be obtained during the offline stage, the channel is approximated by assuming the average channel as the actual channel. The initial sparsification ratio is set to $O^{0}$. The offline stage thus considers the following approximated problem:
@@
where $\gamma^{0}$ is the upper-bound on the initial optimality gap in Theorem 1.
Note that the Problem $\mathcal{P}_{2}$ is still an NP-hard MINLP problem due to the discrete rank value and non-convexity caused by the coupled variables. To address these challenges, we first fix the integer variable $r$ to a specific value and solve the resulting continuous optimization problem $O^{0}$. By iterating over all feasible values of $r$ within the constraint, we can identify the optimal integer $r$. To efficiently solve the continuous optimization problem for a fixed $r$, we present the following lemma, which provides the optimal values of $O^{0}$ given an integer $r$.
@@
$O^{0}=\max\left(O_{min},\min\left(1,\frac{ND_{th}}{AK^{0}r}\right)\right),$ (26)
@@
Consequently, using $O_{min}\!<O^{0}\leq 1$, we can obtain
$O^{0}=\max\left(O_{min},\min\left(1,\frac{ND_{th}}{AK^{0}r}\right)\right),$ (29)
@@
Based on Lemma 2, we can obtain the optimal LoRA rank to be used in training. Subsequently, clients perform fine-tuning based on the specified LoRA rank $r$. At each iteration, an online-stage optimization is conducted to determine the values of $O^{t}$ and $b_{k}^{t}$.